In 2019 Country A had a current account deficit of $1.2 billion. CountryA’s capital account was in a $100 million surplus. In addition, CountryA’s factors of production located in foreign countries earned $600 million.Country A had a trade deficit of $800 million. Assume Country A neithergave nor received unilateral transfers. Country A’s GDP was $9 billion.Answer the following questions about Country A in 2019. and show your work.

(a) What happened to Country A’s net foreign assets during that year?Did it acquire or lose foreign assets?

(b) Compute net factor income from abroad of Country A.

(c) How much income did foreign factors of production earn in CountryA?

(d) Compute Country A’s GNE and GNI.
Answer #1

A. Country's A capital account has a surplus of $100 million. Surplus in capital account means that there is more inflow of foreign exchange as compared to outflows of foreign exchange. This means that Purchase of foreign assets by residents is less. Inflows of foreign currency takes place when there is purchase of assets by foreigners in the domestic country or there is Sale of foreign assets by residents. Surplus in capital account means country A has lost foreign assets.

B. Current account balance = trade balance + balance on account of unilateral transfers + balance on income from investment and work or net factor income from abroad

- $1.2 billion = -$800 million + 0 + net factor income from abroad

-1200, 000, 000 = -800,000,000 + net factor income from abroad

Net factor income from abroad = -1200,000,000 + 800,000,000 = -400,000,000 ( - $ 400 million)

C. Net factor income from abroad = factor income from abroad - factor income paid to abroad

-400million = 600 million - factor income paid to abroad

- 400 million - 600 million = - factor income paid to abroad

Factor Income paid to abroad = 1000 million

Foreign factors of production earned $1000 million from COUNTRY A.

D. GNI = GDP + NET FACTOR INCOME FROM ABROAD

GNI = 9 billion + (-400 million)

GNI = 9000 MILLION - 400 MILLION

GNI = 8600 million

Or 8.6 billion

GDP = GNE + TRADE BALANCE

9 BILLION = GNE + (-800 MILLION)

9 BILLION + 800 MILLION = GNE

GNE = 9.8 billion